Mr. Adams buys a bag of apples and a bag of pears for a total of $14.00. There are eight pounds of fruit in all. The apples cost $1.50 per pound, and the pears cost $2.30 per pound. HOw many pounds of pears did he buy?
Answer by checkley71(8403)

2.30x+1.30(8-x)=14.00
2.30x+10.40-1.30x=14.00
1x=14.00-10.40
x=3.60 pounds of pears.
proof
2.30*3.60+1.30(8-3.60)=14.00
8.28+1.3*4.4=14.00
8.28+5.72=14.00
14.00=14.00